According to ABC Company's standards for variable overhead, each unit produced should require 6 direct labor hours to make. The following information is available for June: 1. actual variable manufacturing overhead cost incurred during June totaled $43,800 2. actual direct labor hours worked during June totaled 8,400 3. the variable overhead spending variance for June was 2,400 favorable 4. the total variable overhead variance for June was 10,140 unfavorable Calculate the number of units produced during June.
Answer #1

Variable overhead efficiency variance = 10140+2400 = 12540 U

Variable overhead spending variance = (Standard rate-actual rate)actual hour

2400 = (X*8400-43800)

-8400X = -46200

X(Standard rate) = 5.50

Variable overhead efficiency variance = (Standard hour-actual hour)Standard rate

-12540 = (X*6*5.5-8400*5.5)

-33X = -33660

X(Unit produced) = 1020 Units